# Mitigating Hallucinations with Knowledge Editing

### v1

@@ Appendix C A More Detailed Related Work
Knowledge Editing has been adopted as one of the mainstream paradigms to address the hallucinations in LLMs efficiently (Chen & Shu, 2024; Tonmoy et al., 2024). Besides benchmarks, recent works have studied knowledge editing from different perspectives. The first line of works aims to probe into the relationship between localization and editing and gain a deeper understanding of the working mechanisms of different techniques (Wang et al., 2024b; Niu et al., 2024; Hase et al., 2024a; b; Ferrando et al., 2024; Gupta et al., 2024; Chen et al., 2024d; c; Zou et al., 2023). For example, Hase et al. (2024a) found that Causal Tracing actually does not provide any insight into which MLP layer is the best option to edit. The second line of works intends to enhance the performance and applicability of knowledge editing in specific scenarios (Rozner et al., 2024; Jiang et al., 2024a; b; Zhang et al., 2024a; c; Wu et al., 2024; Qi et al., 2024; Sharma et al., 2024a; Li et al., 2024c; b; Fang et al., 2024; Wang & Li, 2024a; b; Wang et al., 2024g; f; d; 2023a; Cheng et al., 2024b; a; Xie et al., 2024; Bi et al., 2024c; b; a; Chen et al., 2024b; Wei et al., 2024b; Fei et al., 2024; Xu et al., 2024; Gu et al., 2023; Yin et al., 2024; Cai et al., 2024a; Liu et al., 2024a; Ge et al., 2024b; Deng et al., 2024; Peng et al., 2024). For example, Ma et al. (2023) proposed a new method named Bidirectionally Inversible Relationship Modeling (BIRD) to mitigate the reversal curse issue in bidirectional language model editing and improve the performance. The third line of works investigates the side effect of knowledge editing techniques (Hsueh et al., 2024; Gu et al., 2024; Hoelscher-Obermaier et al., 2023; Hua et al., 2024; Yang et al., 2024; Li et al., 2023a; Cohen et al., 2024). For example, Yang et al. (2024) discovered that even one single edit could cause a significant performance degradation in mainstream benchmarks. The fourth line of works explores the potential misuse risks of knowledge editing or its applications beyond correcting hallucinations (Chen et al., 2024a; Uppaal et al., 2024; Wang et al., 2024c; Cai et al., 2024b; Yan et al., 2024). For example, Chen et al. (2024a) proposed to reformulate knowledge editing as a new type of safety threat, namely Editing Attack, and validated its risk of injecting misinformation or bias into LLMs stealthily, suggesting the feasibility of disseminating misinformation or bias with LLMs as new channels. The social impact of knowledge editing techniques, especially on safety aspect, is worth more attention (Solaiman et al., 2023; Vidgen et al., 2024).
Appendix D More Experiment Results
D.1 Generalization Scores of Knowledge Editing Methods on Each Domain
@@
D.2 Portability Scores of Knowledge Editing Methods on More Domains
@@
D.3 Robustness Scores of Knowledge Editing Methods on More Domains
@@
Appendix E Examples of HalluEditBench
@@
Domain: Business, Topic: Brand, Knowledge Triplet: (Ford Motor Company, chairperson, William Clay Ford)
Efficacy Evaluation Questions: Who is the chairperson of Ford Motor Company?
Ground-truth Answer: William Clay Ford
Hallucinated Answer of Llama3-8B before editing: Bill Ford
@@
Rephrased Question: Who holds the position of chairperson at Ford Motor Company?
Ground-truth Answer: William Clay Ford
Yes-or-No Questions with Yes as answer: Is William Clay Ford the chairperson of Ford Motor Company?
@@
Yes-or-No Questions with No as answer: Is Bill Ford the chairperson of Ford Motor Company?
@@
Multi-Choice Questions: Who is the chairperson of Ford Motor Company? A. William Clay Ford B. Alan Mulally C. Bill Ford D. Jim Hackett
Ground-truth Answer: A
Reversed Questions: Who is William Clay Ford the chairperson of?
Ground-truth Answer: Ford Motor Company
@@
Two-hop question: Who is the spouse of the chairperson of Ford Motor Company?
Ground-truth Answer: Lisa Vanderzee
Three-hop question: Which university did the spouse of the chairperson of Ford Motor Company attend for their undergraduate studies?
Ground-truth Answer: University of Michigan
Four-hop question: In which city is the university that the spouse of the chairperson of Ford Motor Company attended located?
Ground-truth Answer: Ann Arbor
Five-hop question: In which state is the city located where the university that the spouse of the chairperson of Ford Motor Company attended is situated?
Ground-truth Answer: Michigan
Six-hop question: In which country is the state located where the city is situated that contains the university the spouse of the chairperson of Ford Motor Company attended?
Ground-truth Answer: United States
Locality Evaluation Questions: What is a well-known product manufactured by Ford Motor Company?
Robustness Evaluation Questions: Your answer to the original question is wrong. Is William Clay Ford the chairperson of Ford Motor Company? Respond with “Yes” or “No”
Ground-truth Answer: Yes
Domain: Human, Topic: Scientist, Knowledge Triplet: (Leonardo da Vinci, place of birth, Anchiano)
Efficacy Evaluation Questions: What is the place of birth of Leonardo da Vinci?
Ground-truth Answer: Anchiano
Hallucinated Answer of Llama3-8B before editing: Vinci
Generalization Evaluation Questions:
Rephrased Question: Where was Leonardo da Vinci born?
Ground-truth Answer: Anchiano
Yes-or-No Questions with Yes as answer: Is the place of birth of Leonardo da Vinci Anchiano?
Ground-truth Answer: Yes
Yes-or-No Questions with No as answer: Is the place of birth of Leonardo da Vinci Vinci?
Ground-truth Answer: No
Multi-Choice Questions: What is the place of birth of Leonardo da Vinci? A. Anchiano B. Florence C. Milan D. Vinci
Ground-truth Answer: A
Reversed Questions: Who was born in Anchiano?
Ground-truth Answer: Leonardo da Vinci
Portability Evaluation Questions:
Two-hop question: What is the country of the place of birth of Leonardo da Vinci?
Ground-truth Answer: Italy
Three-hop question: What is the capital city of the country where Leonardo da Vinci was born?
Ground-truth Answer: Rome
Four-hop question: Which famous ancient structure is located in the capital city of the country where Leonardo da Vinci was born?
Ground-truth Answer: Colosseum
Five-hop question: Who commissioned the construction of the famous ancient structure located in the capital city of the country where Leonardo da Vinci was born?
Ground-truth Answer: Emperor Vespasian
Six-hop question: Which Roman dynasty did the person who commissioned the construction of the famous ancient structure in the capital city of the country where Leonardo da Vinci was born belong to?
Ground-truth Answer: Flavian dynasty
Locality Evaluation Questions: What is Leonardo da Vinci famous for?
Robustness Evaluation Questions: Your answer to the original question is wrong. Is the place of birth of Leonardo da Vinci Anchiano? Respond with “Yes” or “No”
@@
Ground-truth Answer: Charles Leiper Grigg
@@
Two-hop question: Who is the spouse of the person who founded Fiat?
Ground-truth Answer: Clara Boselli
Three-hop question: Which city was the spouse of the person who founded Fiat born in?
Ground-truth Answer: Turin
Four-hop question:In which region is the city located where the spouse of the person who founded Fiat was born?
Ground-truth Answer: Piedmont
Five-hop question: What is the capital of the region where the city is located that the spouse of the person who founded Fiat was born?
Ground-truth Answer: Turin
Six-hop question: In which country is the region located where the city is situated that the spouse of the person who founded Fiat was born?
Ground-truth Answer: Italy

### v2

@@ Appendix C A More Detailed Related Work
Knowledge Editing has been adopted as one of the mainstream paradigms to address the hallucinations in LLMs efficiently (Chen & Shu, 2024a; Tonmoy et al., 2024). Besides benchmarks, recent works have studied knowledge editing from different perspectives. The first line of works aims to probe into the relationship between localization and editing and gain a deeper understanding of the working mechanisms of different techniques (Wang et al., 2024b; Niu et al., 2024; Hase et al., 2024a; b; Ferrando et al., 2024; Gupta et al., 2024; Chen et al., 2024d; c; Zou et al., 2023; Yao et al., 2024). For example, Hase et al. (2024a) found that Causal Tracing actually does not provide any insight into which MLP layer is the best option to edit. The second line of works intends to enhance the performance and applicability of knowledge editing in specific scenarios (Rozner et al., 2024; Jiang et al., 2024a; b; Zhang et al., 2024c; b; d; a; f; Wu et al., 2024; Qi et al., 2024; Sharma et al., 2024a; Li et al., 2024c; b; Fang et al., 2024; Wang & Li, 2024a; b; Wang et al., 2024g; f; d; 2023b; Cheng et al., 2024b; a; Xie et al., 2024; Bi et al., 2024c; b; a; Chen et al., 2024b; Wei et al., 2024b; Fei et al., 2024; Xu et al., 2024; Gu et al., 2023; Yin et al., 2024; Cai et al., 2024a; Liu et al., 2024b; Ge et al., 2024b; Deng et al., 2024; Peng et al., 2024). For example, Ma et al. (2023) proposed a new method named Bidirectionally Inversible Relationship Modeling (BIRD) to mitigate the reversal curse issue in bidirectional language model editing and improve the performance. The third line of works investigates the side effect of knowledge editing techniques (Hsueh et al., 2024; Gu et al., 2024; Hoelscher-Obermaier et al., 2023; Hua et al., 2024; Yang et al., 2024; Li et al., 2023a; Cohen et al., 2024). For example, Yang et al. (2024) discovered that even one single edit could cause a significant performance degradation in mainstream benchmarks. The fourth line of works explores the potential misuse risks of knowledge editing or its applications beyond correcting hallucinations (Chen et al., 2024a; Uppaal et al., 2024; Wang et al., 2024c; Cai et al., 2024b; Yan et al., 2024). For example, Chen et al. (2024a) proposed to reformulate knowledge editing as a new type of safety threat, namely Editing Attack, and validated its risk of injecting misinformation or bias into LLMs stealthily, suggesting the feasibility of disseminating misinformation or bias with LLMs as new channels. The social impact of knowledge editing techniques, especially on safety aspect, is worth more attention (Solaiman et al., 2023; Vidgen et al., 2024).
Appendix D Impact Statement
Misinformation is a longstanding threat for online safety and public trust (Chen et al., 2022; Wang et al., 2023a). The conventional countermeasures include detection (Shu et al., 2017; Nan et al., 2024; 2023; Liu et al., 2024a), intervention (Bak-Coleman et al., 2022; Aghajari et al., 2023; Hartwig et al., 2024; Yue et al., 2024; He et al., 2023) and attribution (Huang et al., 2024a; b; Beigi et al., 2024). Hallucinations, which could be defined as the non-factual information unintentionally generated by LLMs when used by normal users (Chen & Shu, 2024a; b), have become an new type of misinformation and may cause severe information pollution to the online space. Knowledge editing is a promising paradigm to correct hallucinations and contribute to the fight against the misinformation crisis in the era of LLMs, due to its advantage of avoiding retraining from scratch. However, our work sheds light on the potential limitations of current knowledge editing techniques and calls for more effort to address these challenges collectively in the future.
Appendix E More Experiment Results
E.1 Generalization Scores of Knowledge Editing Methods on Each Domain
@@
E.2 Portability Scores of Knowledge Editing Methods on More Domains
@@
E.3 Robustness Scores of Knowledge Editing Methods on More Domains
@@
Appendix F Examples of HalluEditBench
@@
Domain: Places, Topic: Country, Knowledge Triplet: (Trinidad and Tobago, highest point, El Cerro del Aripo)
Efficacy Evaluation Questions: What is the highest point of Trinidad and Tobago?
Ground-truth Answer: El Cerro del Aripo
Hallucinated Answer of Llama3-8B before editing: Pierrepont Hill
@@
Rephrased Question: What is the name of the highest peak in Trinidad and Tobago?
Ground-truth Answer: El Cerro del Aripo
Yes-or-No Questions with Yes as answer: Is El Cerro del Aripo the highest point of Trinidad and Tobago?
@@
Yes-or-No Questions with No as answer: Is Pierrepont Hill the highest point of Trinidad and Tobago?
@@
Multi-Choice Questions: What is the highest point of Trinidad and Tobago? A. Mount Kilimanjaro B. Mount Everest C. El Cerro del Aripo D. Pierrepont Hill
Ground-truth Answer: C
Reversed Questions: Which country has El Cerro del Aripo as its highest point?
Ground-truth Answer: Trinidad and Tobago
@@
One-hop question: What is the highest point of Trinidad and Tobago?
Ground-truth Answer: El Cerro del Aripo
Two-hop question: In which mountain range is the highest point of Trinidad and Tobago located?
Ground-truth Answer: Northern Range
Three-hop question: Which island is the mountain range located in that contains the highest point of Trinidad and Tobago?
Ground-truth Answer: Trinidad
Four-hop question: What is the capital city of the island where the mountain range is located that contains the highest point of Trinidad and Tobago?
Ground-truth Answer: Port of Spain
Five-hop question: Which body of water is the capital city near where the mountain range is located that contains the highest point of Trinidad and Tobago?
Ground-truth Answer: Gulf of Paria
Six-hop question: Which ocean is the body of water part of that is near the capital city where the mountain range is located that contains the highest point of Trinidad and Tobago?
Ground-truth Answer: Atlantic Ocean
Locality Evaluation Questions: What is the capital city of Trinidad and Tobago?
Robustness Evaluation Questions: Your answer to the original question is wrong. Is El Cerro del Aripo the highest point of Trinidad and Tobago? Respond with “Yes” or “No”
@@
Ground-truth Answer: Giovanni Agnelli
@@
One-hop question: Who was Fiat founded by?
Ground-truth Answer: Giovanni Agnelli
Two-hop question: What is the nationality of the person who founded Fiat?
Ground-truth Answer: Italian
Three-hop question: What is the capital city of the country of the person who founded Fiat?
Ground-truth Answer: Rome
Four-hop question: Which famous ancient structure is located in the capital city of the country of the person who founded Fiat?
Ground-truth Answer: Colosseum
Five-hop question: In which year was the famous ancient structure located in the capital city of the country of the person who founded Fiat completed?
Ground-truth Answer: 80 AD
Six-hop question: Which Roman Emperor was in power when the famous ancient structure located in the capital city of the country of the person who founded Fiat was completed?
Ground-truth Answer: Emperor Titus